OpenStudy (anonymous):

In 2000 the ticket Price was $5.39. The price went to $6.21in 2004. Assuming the linear increase, what would be the approximate price in 2009?? @Algebra

OpenStudy (anonymous):

is it $ 7.235

OpenStudy (anonymous):

The answer would be 7.43 :)

OpenStudy (anonymous):

I agree with sheg. 6.21-5.39 = 0.82 0.82/4 = 0.205 0.205 * 9 = 1.845 5.39+1.845 = $7.235. Or approximately $7.24.